34​% of U.S. adults say they are more likely to make purchases during a sales tax holiday. You randomly select 10 adults. Find the probability that the number of adults who say they are more likely to make purchases during a sales tax holiday is​ (a) exactly​ two, (b) more than​ two, and​ (c) between two and​ five, inclusive.


1
Expert's answer
2021-11-09T09:41:50-0500

Let "X=" the number of adults who say they are more likely to make purchases during a sales tax holiday: "X\sim Bin(n, p)."

Given "n=10, p=0.34, q=1-p=0.66"

(a)


"P(X=2)=\dbinom{10}{2}(0.34)^2(0.66)^{10-2}\approx0.187293"

(b)


"P(X>2)=1-P(X=0)-P(X=1)"

"-P(X=2)=1-\dbinom{10}{0}(0.34)^0(0.66)^{10-0}"

"-\dbinom{10}{1}(0.34)^1(0.66)^{10-1}-\dbinom{10}{2}(0.34)^2(0.66)^{10-2}"

"=1-0.01568336881-0.08079311205"

"-0.18729312338=0.71623039576"

"\approx0.716230"

(c)


"P(2\leq X\leq 5)=P(X=2)+P(X=3)"

"+P(X=4)+P(X=5)"

"=\dbinom{10}{2}(0.34)^2(0.66)^{10-2}"

"+\dbinom{10}{3}(0.34)^3(0.66)^{10-3}"

"+\dbinom{10}{4}(0.34)^4(0.66)^{10-4}"

"+\dbinom{10}{5}(0.34)^5(0.66)^{10-5}"

"\approx0.18729312+0.25729156"

"+0.23195224+0.14338866"

"\approx0.819926"